OpenStudy (epicteatime):

HELP!!! how many moles of aluminum metal would be needed to produce 35.6 grams of Aluminum oxide

OpenStudy (epicteatime):

in the following synthesis reaction... 4 Al + 3 O2 --> 2 Al2O3

OpenStudy (epicteatime):

@ParthKohli

OpenStudy (epicteatime):

@thomaster

OpenStudy (epicteatime):

step by step please

OpenStudy (epicteatime):

...ummm

Parth (parthkohli):

First, how many moles of Al2 O3 is 35.6 grams?

OpenStudy (epicteatime):

sry, im not good with the moles and stuff....

Parth (parthkohli):

OK... what is the molar mass of Al2 O3?

OpenStudy (epicteatime):

101.96g/mol i think

Parth (parthkohli):

OK. No. of moles of Al2 O3 = mass/molar mass = 35.6/101.96.

OpenStudy (epicteatime):

ok

OpenStudy (epicteatime):

then?

Parth (parthkohli):

All right, now let's review the reaction: 4 Al + 3 O2 --> 2 Al2O3 This reaction means that for every 4 moles of Al, 2 moles of Al2 O3 will be formed. Or, for every 1 mole of Al2 O3, there is 2 moles of Al. Right?

OpenStudy (epicteatime):

since you balance it out it becomes that (4 Al + 3 O2 --> 2 Al2O3) right?

Parth (parthkohli):

Yes.

OpenStudy (epicteatime):

ok then what do i do

OpenStudy (epicteatime):

after this one, could you help me with another, little bit more difficult, question???

Parth (parthkohli):

Read the above reply: This reaction means that for every 4 moles of Al, 2 moles of Al2 O3 will be formed. Or, for every 1 mole of Al2 O3, there is 2 moles of Al. Right? We know that there are 2.85 moles of Al2 O3. So the number of moles of Al will be double of 2.85 from what we learned.

OpenStudy (epicteatime):

where did the 2.85 come from

Parth (parthkohli):

oh, sorry... it's 35.6/101.96.

Parth (parthkohli):

I mistakenly calculated 101.96/35.6

OpenStudy (epicteatime):

...?

Parth (parthkohli):

The number of moles of Al2 O3 is 35.6/101.96 = 0.35 moles.

OpenStudy (epicteatime):

yes

Parth (parthkohli):

now, for every one mole of Al2 O3, there's two moles of Al. That means that no. of moles of Al is TWO TIMES no. of moles of Al2 O3.

OpenStudy (epicteatime):

ok

Parth (parthkohli):

So what's the double of 0.35? That's the number of moles of Al which the question requires.

OpenStudy (epicteatime):

0.7?

Parth (parthkohli):

yes.

OpenStudy (epicteatime):

ok, then what?

Parth (parthkohli):

That's all the question requires.

OpenStudy (epicteatime):

what?

OpenStudy (epicteatime):

thats the answer?

OpenStudy (epicteatime):

wow...

Parth (parthkohli):

The question asks for the number of moles of Aluminium. We've calculated the number of moles of Aluminium.

OpenStudy (epicteatime):

ok

OpenStudy (epicteatime):

care to help me with another question?

OpenStudy (epicteatime):

Unbalanced equation: KClO3 g KCl + O2

OpenStudy (epicteatime):

How many moles of oxygen gas can be produced when 78.9 grams of potassium chlorate decompose?

OpenStudy (epicteatime):

i can balance the equation

Parth (parthkohli):

OK, do so.

OpenStudy (epicteatime):

soooo... 2KClO3 --> 2KCl + 3O2?

Parth (parthkohli):

yes, that's right. now - it's given that there are 78.9 grams of potassium chloride. how many moles is that?

OpenStudy (epicteatime):

ummm

OpenStudy (epicteatime):

AHHHH

OpenStudy (epicteatime):

I DONT KNOW!!!

Parth (parthkohli):

Hint: no. of moles = given mass/molar mass

OpenStudy (epicteatime):

uhhhh

OpenStudy (epicteatime):

78.9...

Parth (parthkohli):

given mass = 78.9 g molar mass of KCl = 39 + 35.5 = 74.5

OpenStudy (epicteatime):

ok...

Parth (parthkohli):

divide them to get the number of moles.

OpenStudy (epicteatime):

ok

OpenStudy (epicteatime):

1.0590...

Parth (parthkohli):

Well, that's correct. Now notice that the coefficient of oxygen is 3/2 times that of KCl. So the number of moles of oxygen would also be 3/2 times that of KCl. So what is the number of moles of oxygen then?

OpenStudy (epicteatime):

1.59?

Parth (parthkohli):

Seems good! We're supposed to find the mass of oxygen. Now again use the formula given below: No. of moles = mass/molar mass The molar mass of oxygen is 32 g. No. of moles is 1.59, as you said. Find its mass.

OpenStudy (epicteatime):

20.12?

Parth (parthkohli):

Not quite. 1.59 = mass/32 1.59 * 32 = mass

OpenStudy (epicteatime):

oh ok

OpenStudy (epicteatime):

how do you find the molar mass of oxygen...?

Parth (parthkohli):

The molar mass of O is 16 grams. The molar mass of O2 is 16 * 2 = 32 grams.

OpenStudy (epicteatime):

@whpalmer4

OpenStudy (epicteatime):

@aaronq

OpenStudy (aaronq):

Your question hasn't been answered?

OpenStudy (epicteatime):

yeah, i sorta had to do something else, and didn't get to finish the question...

OpenStudy (aaronq):

what is the question, don't really feel like reading through this whole thread.

OpenStudy (epicteatime):

How many moles of oxygen gas can be produced when 78.9 grams of potassium chlorate decompose? And this is the balanced equation: 2KClO3 --> 2KCl + 3O2

OpenStudy (aaronq):

Okay, so convert the mass of potassium chlorate to moles. -------------------------------------------------------- To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. -------------------------------------------------- Now use those moles to find the moles of \(O_{2~(g)}\) produced using the stoichiometric coefficients. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\)

OpenStudy (epicteatime):

thanks!

OpenStudy (aaronq):

no problem !